# Media Bias Detector: Designing and Implementing a Tool for Real-Time Selection and Framing Bias Analysis in News Coverage

### v1

@@ 1. Introduction
Since the 2016 election, there has been growing concern about the pervasive impact of fake news (Lazer et al., 2018; Vosoughi et al., 2018; Pennycook and Rand, 2021; Allcott and Gentzkow, 2017). However, in practice, most people consume news from mainstream sources, where stories, though factually accurate, can still be biased (Allen et al., 2020). This bias arises from the considerable control that journalists and editors exert over selection (choosing to emphasize some issues, events, or people over others) as well as their framing (choosing the tone, perspective, or facts of a story to present an issue in a particular way) (Entman, 1993; Chong and Druckman, 2007; Rothschild et al., 2023). Unlike outright falsehoods, selection and framing can subtly mislead readers, making these forms of bias difficult for traditional fact-checking tools to detect (Park et al., 2021). Experimental evidence has shown that it is remarkably easy to mislead a reader without making any explicitly false statements (Alexander, 2022; Lifchits et al., 2021). Moreover, because factually accurate but biased news coverage tends to be consumed by much larger populations than categorically fake news, its negative impact can be considerably larger in aggregate (Allen et al., 2024).
Measuring and exposing media bias, which we define as the preferential selection of some stories, facts, people, events, or perspectives over others, is critical to challenging the illusion of objectivity promoted by many news organizations. By revealing how editorial choices (decisions about which facts, voices, and perspectives to emphasize or downplay) shape coverage, readers can better understand both the context of the information they consume and the perspectives they are missing. A lack of diverse perspectives can contribute to echo chambers, where individuals relying on a single source or a limited set of ideologically similar sources are more likely to encounter “separate realities,” fostering disagreement on basic facts or priorities (Muise et al., 2022; Hosseinmardi et al., 2023; Bram, 2024).
@@
Communications researchers have long emphasized the importance of media bias (Stroud, 2008; Baum and Groeling, 2008; Chong and Druckman, 2007; Xiang and Sarvary, 2007; Chiang and Knight, 2011); however, methodological limitations have until recently prevented its identification at scale. Traditional approaches, like manually quantifying topics, viewpoints, or facts across publishers, are costly and time-consuming (Budak et al., 2016; Lim et al., 2020; Mitchell et al., 2017), while natural language processing (NLP) methods such as n-gram counting typically fail to capture the full context of an article (Gentzkow and Shapiro, 2010; D’Alessio and Allen, 2000). Now, with advancements in large language models (LLMs), it has become possible to annotate large documents faster and more efficiently, while maintaining high accuracy (Goel et al., 2023). Specifically, LLMs can efficiently ingest and parse tens of thousands of articles, extracting key features such as topics, subtopics, and facts as well as labeling them according to subjective quantities such as tone (positive vs. negative) and partisanship (Democrat vs. Republican) (Peña et al., 2023; Xu et al., 2023). This ability to extract nuanced information from large datasets makes LLMs a valuable tool for identifying and analyzing media bias.
In this paper, we present the Media Bias Detector, an LLM-driven tool designed to dynamically track and analyze major news coverage from a diverse collection of prominent publishers across the political spectrum. [^1] We limit our initial investigation to ten mainstream news publishers to maintain data quality and keep labeling costs manageable. But, the tool is easily adaptable to additional publishers. We operationalize media bias in two distinct ways, corresponding to the aforementioned distinction between selection and framing bias. First, we quantify selection bias by measuring the differential attention paid to different news categories (e.g. politics vs. business) as well as topics (e.g. the 2024 election within politics) and subtopics (e.g. the presidential horse race within the 2024 election). Second, we quantify framing bias along two key dimensions: a) political lean, which we define as the extent to which an article aligns with the viewpoints, policies, or concerns of Republicans versus Democrats, either explicitly or implicitly; and b) tone, defined as the emotional slant of coverage (i.e. positive, neutral, negative). Unlike existing tools, which typically consider subsets of news stories and label publishers broadly as left, right, or center, the Media Bias Detector annotates individual articles produced every day. These annotations are then aggregated to provide a dynamic, data-driven view of each publisher, reflecting the diversity of their coverage rather than assigning static ideological labels. Our inclusion of tone also addresses a gap in media bias detection, which often focuses on political lean but has paid less attention to sentiment as a source of bias (Robertson et al., 2023). This added granularity allows readers to explore the distribution of lean and tone across articles published by a news outlet. For instance, a user might find that while the Wall Street Journal appears generally neutral overall, its coverage leans more Republican on immigration and slightly more Democratic on the environment and reproductive rights.
@@ 2. Background and Related Work
Bias, in statistical theory, implies the presence of a ”true” value from which the biased estimate systematically differs. However, in practice, we generally lack a ground truth, making it infeasible to directly assess bias for individual articles or even entire publishers. This assertion might seem counterintuitive, as we often perceive articles as biased when we read them. This perception arises because we implicitly treat our subjective opinions as the ”truth,” even though they are likely biased as well (Pronin et al., 2004). In the absence of a universally accepted ground truth, we instead need to identify bias by observing patterns and differences in how news is reported between publishers, or within publishers over time. In this section, we anchor our discussion of media bias in the concepts of selection and framing biases and consider their relevance to designing media bias tools grounded in HCI principles. We also situate the Media Bias Detector among existing media bias detection tools and discuss their limitations. Finally, we review current literature on LLMs for annotation tasks and explain why these models are well-suited for analyzing media bias.
@@ 2.1. Agenda-Setting, Framing, and the Role of HCI in Designing Tools for Media Bias
Media bias presents major implications for the HCI community as online platforms increasingly become the primary medium for news consumption. Within the communications field, media bias is often analyzed through two key theories: agenda-setting and framing. Agenda-setting theory argues that the media doesn’t tell us what to think but rather what to think about (McCombs and Shaw, 1972, 2005). The process of selection bias—choosing certain stories or facts to report while omitting others—directly contributes to the media’s agenda-setting power, shaping public discourse by directing attention to specific topics over others (Dearing and Rogers, 1996). For example, a news outlet might publish many stories on a frontrunner’s campaign rallies while giving limited coverage to other candidates. Agenda-setting theory is highly relevant to HCI because digital platforms increasingly mediate what content users see (Naser, 2020). Algorithms, interface design, and personalized recommendation systems often act as agenda-setters, determining what content is presented or excluded from users’ screens (Maudet, 2019; Eslami et al., 2016). This phenomenon raises critical questions in HCI about how systems shape user awareness and understanding, as well as how tools can be designed to challenge these embedded biases.
@@
Beyond selecting which issues to emphasize, the media also influences public perception via the manner in which information is presented—a process known as framing (Weaver, 2007). Framing bias builds on selection bias by not just determining what is covered, but how it is covered. This involves the inclusion or omission of specific details and perspectives, the tone or language used, the context provided, or the omission of background information, all of which can significantly alter how an issue is perceived by the audience (Entman, 1993; Gentzkow and Shapiro, 2006; Chong and Druckman, 2007). For instance, framing a protest as a ”riot” in one article versus a ”peaceful demonstration” in another can significantly influence how readers perceive the event (Brown and Mourão, 2021; Susánszky et al., 2022). Similarly, framing immigration as an ”opportunity for economic growth” versus an ”immigration crisis” can sway public perception on the issue (Igartua and Cheng, 2009). Framing effects relate to HCI through a shared focus on how the presentation and emphasis of information influences its importance or impact (Hartmann et al., 2008). Media bias tools informed by HCI principles can play a critical role in revealing framing biases (Dingler et al., 2020). For example, tools that allow users to analyze data from multiple perspectives (Ainsworth, 2008), explore specific points in more detail (Springer and Whittaker, 2019), or compare diverse sources (Bhuiyan et al., 2023; Bråten and Strømsø, 2011) can foster a deeper understanding of how the same event is portrayed in different ways.
@@
Although the concepts of selection and framing bias are well-known in media studies and communications (O’neill and Harcup, 2009; D’Alessio and Allen, 2000; Matthes, 2009), their impact has become even more pronounced in today’s digital age (Harcup and O’neill, 2017; Bourgeois et al., 2018). The rapid dissemination and broad reach of online news allow information consumers to quickly access content that reinforces their pre-existing opinions. Research indicates that online news consumption exhibits a polarized pattern, with users spending significantly more time on news sources that align with their political leanings compared to those that do not (Garimella et al., 2021). Tools that make media bias visible—especially in terms of selection and framing biases—can help users recognize these patterns and engage more critically with their news consumption (Kiss et al., 2023). By applying HCI principles to these tools, we can design interfaces that promote media literacy and encourage balanced engagement with digital news platforms.
@@ 2.2. Examining Current Approaches and Limitations in Assessing Media Bias
Scholars have proposed numerous taxonomies to understand media bias, yet no universally accepted set of media bias metrics or standard measurements exist (Puglisi and Snyder Jr, 2015; Gentzkow et al., 2015; Morstatter et al., 2018; Hamborg et al., 2019; Huang et al., 2024). Current methods often reduce news publishers to a single metric by assigning a political lean rating from ”Left” to ”Right” (AllSides, 2024; Check, 2024; News, 2024). These tools simplify bias labels to make it easier for users to digest, such as using 2-dimensional axes with political lean and reliability (Media, nd) or a 1-dimensional categorization of publishers across the political spectrum (AllSides, 2024). Although useful, these static classifications fail to capture within-publisher differences in bias over time, across topics, or even across individual articles.
Evaluating bias at the article level is less common and more challenging. Earlier tools evaluating article level bias relied on basic NLP techniques, such as keyword frequency analysis, to cluster similar stories in the media (Park et al., 2009). However, these methods are limited in their ability to account for context and the evolving nature of news stories. Recent advancements in LLMs offer improved contextual understanding capabilities which can more accurately help identify media bias in news coverage. Even though other proprietary tools that incorporate AI features have emerged (Biasly, 2024; AllSides, 2024), these platforms generally do not update to reflect media bias in recent news coverage. Furthermore, the lack of transparency around their AI systems makes it unclear which models they use and how they arrive at their conclusions.
@@ 2.3. Incorporating Large Language Models (LLMs) for Media Bias Detection
LLMs, such as GPT-4, utilize deep contextual embeddings to capture subtle semantic relationships and nuances within text (Brown, 2020; Achiam et al., 2023; Radford et al., 2019). Studies have shown that LLMs perform well in generative tasks such as summarization (Liu et al., 2024; Ravaut et al., 2024; Liu et al., 2023), as well as discriminative tasks like sentiment analysis and topic classification (Chang et al., 2024; Peña et al., 2023; Zhang et al., 2023a). More recent studies have shown that state-of-the-art LLMs exhibit increasingly complex reasoning and problem-solving capabilities (Bubeck et al., 2023). These findings suggest that LLMs can be especially useful for the complex task of media bias detection, where understanding context, tone, and subtle language is crucial.
@@
While traditional machine learning models used in annotation tasks require task-specific training, instruction-tuned pre-trained LLMs can generalize to different tasks in a zero-shot setting (Wang et al., 2024). Prior work has shown that using LLMs to automate the initial stages of data annotation enables researchers to quickly process large volumes of content that would be unmanageable with manual annotation alone (Ziems et al., 2024). This scalability is critical given the speed with which major news outlets publish new articles, which must be analyzed in near real-time to capture evolving narratives and help readers interpret the news. By incorporating LLMs into the Media Bias Detector’s news analysis pipeline, we can efficiently extract detailed information like sentence-level codings and classifications by topic, subtopic, article type, tone, and political lean, while significantly increasing the volume of articles we can annotate. However, it is important to acknowledge potential limitations of LLMs in zero-shot settings, such as political biases in their outputs and the influence of this bias on the tool’s results. These challenges are explored in greater depth in Section 9.
@@
The integration of LLMs into traditionally human processes is informed by a growing body of research that shows that human-AI collaboration enhances the accuracy and reliability of automated systems (Uchendu et al., 2023; Goel et al., 2023). Although LLMs can quickly process large datasets and perform initial classifications, human oversight is crucial for providing context that LLMs do not have for current events and verifying the model’s outputs for consistent standards of accuracy (Wang et al., 2024; Amirizaniani et al., 2024). In the Media Bias Detector, humans play an important role at every step, from generating a targeted list of relevant news topics for LLMs to classify, to continuously monitoring the model’s classifications over time.
@@
We intentionally chose to evaluate the responses generated by LLMs through a process of human annotation that emphasizes validation over independent labeling. Rather than having annotators blindly label the data, we engage them in a validation task where they read GPT-generated responses to assess their coherence and soundness. Although having humans independently label content and then comparing it to GPT’s output could yield interesting insights, it is important to recognize that human disagreements often occur, even among reasonable individuals. Particularly in complex tasks, such as reading an entire article and labeling its lean and tone, as well as extracting facts, achieving high agreement is challenging (Mitchell et al., 2018). Our methodology accounts for this inherent subjectivity and aims to ensure that topics presented on the Media Bias Detector are coherent, reasonable, and free from overt inaccuracies. By involving humans in the validation process, we impose a layer of quality control that acknowledges the absence of a singular ground truth.
@@ 3. Design Considerations
Media bias is a multidimensional problem (Eberl et al., 2017) and cannot accurately be captured by just one metric. A central goal of our tool is to support breadth of exploration, providing users with affordances that enable them to examine the data from multiple perspectives and draw their own conclusions. This design is informed by principles of learning with multiple representations, which enhance cognitive processes in learning by allowing users to approach complex data from different angles (Ainsworth, 2008, 1999). To operationalize these principles, we use faceted categories to structure the interface, enabling users to dynamically toggle between variables and uncover relationships that might otherwise be difficult to identify (Yee et al., 2003; Lee et al., 2009). For example, users can switch between metrics such as coverage volume, political lean, and tone across publishers and topics. They can also adjust the date range to compare publishers over any time frame and filter by article type, allowing for a comprehensive comparison of how various content categories reflect media bias over selected time periods. Figures 1 and 2 showcase how the Media Bias Detector incorporates faceted categories to help users process information through multiple representations.
@@
In addition to being a multidimensional problem, media bias is also a multilevel one. While overall news coverage on a particular category may not differ much between two publishers, how they choose to cover specific topics, subtopics, and events within that category can reveal a hidden bias. A major affordance of our tool is the ability to engage in deep exploration of the data after seeing an overview, allowing users to uncover more detailed biases without being overwhelmed at the very first stage. This approach, informed by the principle of progressive disclosure, affords a gradual reveal of complexity as users build familiarity with the system (Carroll and Carrithers, 1984; Smith et al., 1982; Nakatani and Rohrlich, 1983; Springer and Whittaker, 2019). From a media bias standpoint, we apply progressive disclosure by allowing users to begin with broad categories and gradually narrow their focus to more specific areas. As users engage with a particular category, they can drill down into increasingly detailed topics and subtopics, enabling a deeper exploration of media coverage. This also extends beyond topic-level analysis, as our tool affords sentence-level text annotation which helps users examine how different news publishers write about the same events. Users can begin by exploring a news event cluster, dive into the top facts reported about that event, and then compare how the same facts are presented across different publications. Figure 3 illustrates how a user can progressively refine their analysis at the category, topic, and subtopic levels.
@@
News readers often have limited time and tend to rely on a single source or ideologically aligned sources to learn about current events (Iyengar and Hahn, 2009). The affordance of comparative analysis within our tool allows users to easily identify similarities and differences between multiple publishers in one place, which may help them recognize discrepancies between perceived and actual coverage (Bhuiyan et al., 2023; Stahl et al., 1996; Bråten and Strømsø, 2011). To reduce cognitive load, our tool provides side-by-side comparisons of publishers and topics with the goal of better contextualizing the broader news ecosystem without users needing to switch between sources.

### v2

@@ 1. Introduction
[…]
@@
Communications researchers have long emphasized the importance of media bias (stroud2008media; baum2008new; chong2007framing; xiang2007news; chiang2011media); however, methodological limitations have until recently prevented its identification at scale. Traditional approaches, like manually quantifying topics, viewpoints, or facts across publishers, are costly and time-consuming (budak2016fair; lim2020annotating; mitchell2017covering), while natural language processing (NLP) methods such as n-gram counting typically fail to capture the full context of an article (gentzkow2010drives; d2000media). Now, with advancements in large language models (LLMs), it has become possible to annotate large documents faster and more efficiently, while maintaining high accuracy (goel2023llms). Specifically, LLMs can efficiently ingest and parse tens of thousands of articles, extracting key features such as topics, subtopics, and facts as well as labeling them according to subjective quantities such as tone (positive vs. negative) and partisanship (Democrat vs. Republican) (pena2023leveraging; xu2023large). This ability to extract nuanced information from large datasets makes LLMs a valuable tool for identifying and analyzing media bias.
[…]
@@ 2. Background and Related Work
Bias, in statistical theory, implies the presence of a ”true” value from which the biased estimate systematically differs. However, in practice, we generally lack a ground truth, making it infeasible to directly assess bias for individual articles or even entire publishers. This assertion might seem counterintuitive, as we often perceive articles as biased when we read them. This perception arises because we implicitly treat our subjective opinions as the ”truth,” even though they are likely biased as well (pronin2004objectivity). In the absence of a universally accepted ground truth, we instead need to identify bias by observing patterns and differences in how news is reported between publishers, or within publishers over time. In this section, we anchor our discussion of media bias in the concepts of selection and framing biases and consider their relevance to designing media bias tools grounded in HCI principles. We also situate the Media Bias Detector among existing media bias detection tools and discuss their limitations. Finally, we review current literature on LLMs for annotation tasks and explain why these models are well-suited for analyzing media bias.
@@ 2.1. Agenda-Setting, Framing, and the Role of HCI in Designing Tools for Media Bias
Media bias presents major implications for the HCI community as online platforms increasingly become the primary medium for news consumption. Within the communications field, media bias is often analyzed through two key theories: agenda-setting and framing. Agenda-setting theory argues that the media doesn’t tell us what to think but rather what to think about (agenda_setting_1972; mccombs2005agenda). The process of selection bias—choosing certain stories or facts to report while omitting others—directly contributes to the media’s agenda-setting power, shaping public discourse by directing attention to specific topics over others (dearing1996agenda). For example, a news outlet might publish many stories on a frontrunner’s campaign rallies while giving limited coverage to other candidates. Agenda-setting theory is highly relevant to HCI because digital platforms increasingly mediate what content users see (Naser2020RelevanceAC). Algorithms, interface design, and personalized recommendation systems often act as agenda-setters, determining what content is presented or excluded from users’ screens (maudet; Eslami2016FirstI). This phenomenon raises critical questions in HCI about how systems shape user awareness and understanding, as well as how tools can be designed to challenge these embedded biases.
@@
Beyond selecting which issues to emphasize, the media also influences public perception via the manner in which information is presented—a process known as framing (weaver2007thoughts). Framing bias builds on selection bias by not just determining what is covered, but how it is covered. This involves the inclusion or omission of specific details and perspectives, the tone or language used, the context provided, or the omission of background information, all of which can significantly alter how an issue is perceived by the audience (entman1993framing; gentzkow2006media; chong2007framing). For instance, framing a protest as a ”riot” in one article versus a ”peaceful demonstration” in another can significantly influence how readers perceive the event (brown_protest_2021; susanszky2022media). Similarly, framing immigration as an ”opportunity for economic growth” versus an ”immigration crisis” can sway public perception on the issue (Igartua2009ModeratingEO). Framing effects relate to HCI through a shared focus on how the presentation and emphasis of information influences its importance or impact (hartmann_framing). Media bias tools informed by HCI principles can play a critical role in revealing framing biases (dingler_workshop). For example, tools that allow users to analyze data from multiple perspectives (ainsworth2008educational), explore specific points in more detail (springer2019progressive), or compare diverse sources (bhuiyan2023newscomp; braaten2011measuring) can foster a deeper understanding of how the same event is portrayed in different ways.
@@
Although the concepts of selection and framing bias are well-known in media studies and communications (o2009news; d2000media; matthes2009s), their impact has become even more pronounced in today’s digital age (harcup2017news; bourgeois2018selection). The rapid dissemination and broad reach of online news allow information consumers to quickly access content that reinforces their pre-existing opinions. Research indicates that online news consumption exhibits a polarized pattern, with users spending significantly more time on news sources that align with their political leanings compared to those that do not (garimella2021political). Tools that make media bias visible—especially in terms of selection and framing biases—can help users recognize these patterns and engage more critically with their news consumption (markus). By applying HCI principles to these tools, we can design interfaces that promote media literacy and encourage balanced engagement with digital news platforms.
@@ 2.2. Examining Current Approaches and Limitations in Assessing Media Bias
Scholars have proposed numerous taxonomies to understand media bias, yet no universally accepted set of media bias metrics or standard measurements exist (puglisi2015empirical; gentzkow2015media; morstatter_2018; hamborg2019automated; huang2024uncovering). Current methods often reduce news publishers to a single metric by assigning a political lean rating from ”Left” to ”Right” (allsides; mediabiasfactcheck; groundnews). These tools simplify bias labels to make it easier for users to digest, such as using 2-dimensional axes with political lean and reliability (adfontesmedia) or a 1-dimensional categorization of publishers across the political spectrum (allsides). Although useful, these static classifications fail to capture within-publisher differences in bias over time, across topics, or even across individual articles.
Evaluating bias at the article level is less common and more challenging. Earlier tools evaluating article level bias relied on basic NLP techniques, such as keyword frequency analysis, to cluster similar stories in the media (park2009newscube). However, these methods are limited in their ability to account for context and the evolving nature of news stories. Recent advancements in LLMs offer improved contextual understanding capabilities which can more accurately help identify media bias in news coverage. Even though other proprietary tools that incorporate AI features have emerged (biasly; allsides), these platforms generally do not update to reflect media bias in recent news coverage. Furthermore, the lack of transparency around their AI systems makes it unclear which models they use and how they arrive at their conclusions.
@@ 2.3. Incorporating Large Language Models (LLMs) for Media Bias Detection
LLMs, such as GPT-4, utilize deep contextual embeddings to capture subtle semantic relationships and nuances within text (brown2020language; achiam2023gpt; radford2019language). Studies have shown that LLMs perform well in generative tasks such as summarization (liu-etal-2024-learning; ravaut-etal-2024-context; liu2023learning), as well as discriminative tasks like sentiment analysis and topic classification (chang2024survey; pena2023leveraging; zhang2023sentiment). More recent studies have shown that state-of-the-art LLMs exhibit increasingly complex reasoning and problem-solving capabilities (bubeck2023sparks). These findings suggest that LLMs can be especially useful for the complex task of media bias detection, where understanding context, tone, and subtle language is crucial.
@@
While traditional machine learning models used in annotation tasks require task-specific training, instruction-tuned pre-trained LLMs can generalize to different tasks in a zero-shot setting (wang_humanllm_2024). Prior work has shown that using LLMs to automate the initial stages of data annotation enables researchers to quickly process large volumes of content that would be unmanageable with manual annotation alone (ziems2024can). This scalability is critical given the speed with which major news outlets publish new articles, which must be analyzed in near real-time to capture evolving narratives and help readers interpret the news. By incorporating LLMs into the Media Bias Detector’s news analysis pipeline, we can efficiently extract detailed information like sentence-level codings and classifications by topic, subtopic, article type, tone, and political lean, while significantly increasing the volume of articles we can annotate. However, it is important to acknowledge potential limitations of LLMs in zero-shot settings, such as political biases in their outputs and the influence of this bias on the tool’s results. These challenges are explored in greater depth in Section 9.
@@
The integration of LLMs into traditionally human processes is informed by a growing body of research that shows that human-AI collaboration enhances the accuracy and reliability of automated systems (uchendu2023does; goel2023llms). Although LLMs can quickly process large datasets and perform initial classifications, human oversight is crucial for providing context that LLMs do not have for current events and verifying the model’s outputs for consistent standards of accuracy (wang_humanllm_2024; amirizaniani2024developing). In the Media Bias Detector, humans play an important role at every step, from generating a targeted list of relevant news topics for LLMs to classify, to continuously monitoring the model’s classifications over time.
@@
We intentionally chose to evaluate the responses generated by LLMs through a process of human annotation that emphasizes validation over independent labeling. Rather than having annotators blindly label the data, we engage them in a validation task where they read GPT-generated responses to assess their coherence and soundness. Although having humans independently label content and then comparing it to GPT’s output could yield interesting insights, it is important to recognize that human disagreements often occur, even among reasonable individuals. Particularly in complex tasks, such as reading an entire article and labeling its lean and tone, as well as extracting facts, achieving high agreement is challenging (mitchell2018distinguishing). Our methodology accounts for this inherent subjectivity and aims to ensure that topics presented on the Media Bias Detector are coherent, reasonable, and free from overt inaccuracies. By involving humans in the validation process, we impose a layer of quality control that acknowledges the absence of a singular ground truth.
@@ 3. Design Considerations
Media bias is a multidimensional problem (eberl2017one) and cannot accurately be captured by just one metric. A central goal of our tool is to support breadth of exploration, providing users with affordances that enable them to examine the data from multiple perspectives and draw their own conclusions. This design is informed by principles of learning with multiple representations, which enhance cognitive processes in learning by allowing users to approach complex data from different angles (ainsworth2008educational; ainsworth1999functions). To operationalize these principles, we use faceted categories to structure the interface, enabling users to dynamically toggle between variables and uncover relationships that might otherwise be difficult to identify (hearst_faceted; lee_faceted). For example, users can switch between metrics such as coverage volume, political lean, and tone across publishers and topics. They can also adjust the date range to compare publishers over any time frame and filter by article type, allowing for a comprehensive comparison of how various content categories reflect media bias over selected time periods. Figures 1 and 2 showcase how the Media Bias Detector incorporates faceted categories to help users process information through multiple representations.
@@
In addition to being a multidimensional problem, media bias is also a multilevel one. While overall news coverage on a particular category may not differ much between two publishers, how they choose to cover specific topics, subtopics, and events within that category can reveal a hidden bias. A major affordance of our tool is the ability to engage in deep exploration of the data after seeing an overview, allowing users to uncover more detailed biases without being overwhelmed at the very first stage. This approach, informed by the principle of progressive disclosure, affords a gradual reveal of complexity as users build familiarity with the system (carroll_1984; smith1982star; nakatani1983soft; springer2019progressive). From a media bias standpoint, we apply progressive disclosure by allowing users to begin with broad categories and gradually narrow their focus to more specific areas. As users engage with a particular category, they can drill down into increasingly detailed topics and subtopics, enabling a deeper exploration of media coverage. This also extends beyond topic-level analysis, as our tool affords sentence-level text annotation which helps users examine how different news publishers write about the same events. Users can begin by exploring a news event cluster, dive into the top facts reported about that event, and then compare how the same facts are presented across different publications. Figure 3 illustrates how a user can progressively refine their analysis at the category, topic, and subtopic levels.
@@
News readers often have limited time and tend to rely on a single source or ideologically aligned sources to learn about current events (iyengar2009red). The affordance of comparative analysis within our tool allows users to easily identify similarities and differences between multiple publishers in one place, which may help them recognize discrepancies between perceived and actual coverage (bhuiyan2023newscomp; stahl1996happens; braaten2011measuring). To reduce cognitive load, our tool provides side-by-side comparisons of publishers and topics with the goal of better contextualizing the broader news ecosystem without users needing to switch between sources.